OpenStudy (anonymous):

A tile factory earns money by a flat fee for delivery and a sales price per tile sold. For one customer, the tile factory shipped 10,000 units to a retail store at a rate of $0.25 per tile. The total the tile factory earned from that sale was $3,000. Which of the following equations describes the revenue of the tile factory for this sale in terms of tiles sold? A. y = 0.25x + 10,000 B. y = 0.25x + 3,000 C. y = 0.25x + 500 D. y = 0.25x + 0.25

undeadknight26 (undeadknight26):

well since its .25 per tile it would be .25x + 3000 i believe...

undeadknight26 (undeadknight26):

sorry in advance if i am wrong.

OpenStudy (anonymous):

thats what im thinking but i think thats the total you get from all it so why put it as the answer

OpenStudy (anonymous):

"10,000 units to a retail store at a rate of $0.25 per tile. The total the tile factory earned from that sale was $3,000." factory earned: $3,000 tiles sold 10,000 price per tile: $0.25 3,000 = 0.25*10,000 + b solve for b b=500 so, 3,000 = 0.25*10000+500 or y = 0.25x + 500

OpenStudy (anonymous):

yes thats the answer i will check right now i was debating bettween b and c